OpenStudy (anonymous):

4x+7y=6 What is the slope of the line given by the linear equation above?

OpenStudy (cwrw238):

7y = -4x + 6 slope = -4/7

OpenStudy (anonymous):

make it into slope intercept form. so 7y = -4x + 6. now isolate Y and make it one, so divide by 7. So y = -4/7x + 6. y = mx+b, so "m" or the slope is - 4/7x.

OpenStudy (anonymous):

oops didn't see that he already answered. oh well lol